A coffee merchant has coffee beens that sell for $9 p/lb & $12 p/lb. the two types are mixed to to create 100 lb of a mixture to sell for $11.25 p/lb. How much of each type of bean should be used in the mixture?
Answer by checkley75(3666)

100*11.25=x*12+(100-x)9
1125=12x+900-9x
12x-9x=1125-900
3x=225
x=225/3
x=75 pounds of $12 coffee.
100-75=25 pounds of $9 coffee.
proof
100*11.25=75*12+25*9
1125=900+225
1125=1125